A seed company sells two grades of grass seed. A 100-pound bag of a mixture of rye and KY bludgrass sells for $245 and a 100 pound bag of bluegrass sells for $347. How many bags of each are sold in a week when the receipts for 19 bags are $5,369.
I believe the answer is 12 bags of the one priced $245 and 7 bags that sell for $347. However, I doubt that I solved the problem in the correct manner. Will you please tell me the proper way to solve such a problem?
Thank you.
ISBN: 0-534-49394-7
Intermediate Algebra (Alan S. tussy and R. David Gustafson (Third Edition)

Answer by ptaylor(2198)

GOOD WORK!!!! I DON'T KNOW HOW YOU SOLVED THE PROBLEM, BUT I THINK THAT YOU GOT THE CORRECT ANSWER---------AT LEAST, WE GOT THE SAME ANSWER.

Let x=number of bags of the $245 brand that are sold
Then 19-x =number of bags of the $347 brand that are sold
Now we know that:
receipts for the $245 brand are 245x; receipts for the $347 brand are 347(19-x) and we are told that these add up to $5369 so our equation to solve is:
245x+347(19-x)=5369 get rid of parens
245x+6593-347x=5369 subtract 6593 fron both sides
245x-347x=5369-6593 collect like terms
-102x=-1224 divide both sides by -102
x=12 bags of the $245 brand
19-x=19-12=7 bags of the $347 brand

Hope this helps---ptaylor